$SPAD/src/input drawx.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
)clear all
--Plotting Two Dimensional Functions of One Variable
draw(sin(tan(x))-tan(sin(x)),x=0..6)
draw(sin(tan(x))-tan(sin(x)),x=0..6,unit == [1.0,0.5],toScale == true)
--Plotting Two Dimensional Parametric Plane Curves
draw(curve(9*sin(3*t/4),8*sin(t)),t=-4*%pi..4*%pi)
draw(curve(9*sin(3*t/4),8*sin(t)),t=-4*%pi..4*%pi,coordinates == polar)
--Plotting Three Dimensional Parametric Space Curves
draw(curve(cos(t),sin(t),t),t=0..6)
draw(curve(cos(t),sin(t),t),t=0..6,tubeRadius == .35,tubePoints == 8)
--Plotting Three Dimensional Functions of Two Variables
draw(cos(x*y),x = -3..3, y = -3..3)
l(x:DoubleFloat,y:DoubleFloat):DoubleFloat == cos(x*y)
colorFxn(x:DoubleFloat,y:DoubleFloat):DoubleFloat == 1/(x**2 + y**2 + 1)
draw(cos(x*y),x = -3..3, y = -3..3, colorFunction == colorFxn)
--Plotting Three Dimensional Parametric Surfaces
draw(surface(u*cos(v),u*sin(v),v*cos(u)),u=-4..4,v=0..2*%pi)
l(x:DoubleFloat,y:DoubleFloat):DoubleFloat == cos(x*y)
colorFxn(x:DoubleFloat,y:DoubleFloat):DoubleFloat == 1/(x**2 + y**2 + 1)
draw(cos(x*y),x = -3..3, y = -3..3, colorFunction == colorFxn)

———-

References^{[1] nothing}

4